# Preference-Based Gradient Estimation for ML-Based Approximate Combinatorial Optimization

### v1

@@ B.2 Probability of Finding the Optimal TSP-Tour
In other words, in the limit, the insertion algorithm steered by $\bm{s}$ always finds optimal solution $\bm{y}$.
Proof.
We know that the solution produced by any insertion algorithm $h(\hat{\bm{y}}\mid{\mathcal{G}})$ is within $\lceil\log(|V|)\rceil+1$ of the optimal solution (Rosenkrantz et al., 1977), i.e.
$J(\hat{\bm{y}})\leq\big{(}\lceil\log(|V|)\rceil+1\big{)}\>J(\bm{y})\qquad\textup{for }\hat{\bm{y}}\sim h(\hat{\bm{y}}\mid{\mathcal{G}}).$
Let ${\mathcal{G}}^{\prime}$ be the graph that the insertion algorithm receives as input after modification through $\bm{s}$, i.e. $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})=h(\hat{\bm{y}}\mid{\mathcal{G}}^{\prime})$. As $\sigma(\bm{s})$ approaches $\bm{y}$, the weights in ${\mathcal{G}}^{\prime}$ of edges in $\bm{y}$ approach 0. Meanwhile, the weights of the other edges approach their original weights from ${\mathcal{G}}$, which are greater than 0. This means that the length of the tour $\bm{y}$ in ${\mathcal{G}}^{\prime}$ approaches 0, while the lengths of any other tour in ${\mathcal{G}}^{\prime}$ approach values greater than 0.
This means that the maximum length of tours that the insertion algorithm can find approaches 0, and, as soon as the lengths of the suboptimal tours in ${\mathcal{G}}^{\prime}$ are far enough away from 0, $\bm{y}$ is the only tour left that the insertion algorithm can find. ∎
Appendix C Details Regarding Baselines
C.1 Supervised Training with Binary Cross Entropy Loss
The task is treated as an edge-level binary classification task. The network is trained using a binary cross entropy loss:
@@ C.1 Supervised Training with Binary Cross Entropy Loss
C.2 REINFORCE
Recall that the REINFORCE algorithm, also known as the score function estimator, calculates
@@ C.2 REINFORCE
  $\nabla_{\theta}J(y)\leftarrow J(y)\left(y-\displaystyle\frac{1}{N}\sum_{i=1}^{N}y_{i}\right)$               where $\varepsilon_{i}\sim\mathrm{Gumbel}(0,1)$                           $y_{i}\leftarrow h(\theta+\varepsilon_{i})$
@@
C.3 Implicit Maximum Likelihood Estimator (I-MLE)
@@
In the backward pass, assume we know the gradient of the loss w.r.t. to $\bm{z}$, $\nabla_{\bm{z}}\mathcal{L}$. First, obtain a modified input $\bm{\theta}^{\prime}$ for which we can expect better outputs compared to $\bm{\theta}$. One generally applicable option suggested by (Niepert et al., 2021) is $\bm{\theta}^{\prime}=\bm{\theta}-\lambda\nabla_{\bm{z}}\mathcal{L}$, where $\lambda$ is a hyperparameter. Using the same noise $\bm{\epsilon}$ as in the forward pass, perturb $\bm{\theta}^{\prime}$ and run the CO solver on $\bm{\theta}^{\prime}+\bm{\epsilon}$, obtaining $\bm{z}^{\prime}$. Finally, return the estimated gradient $\nabla_{\bm{\theta}}\mathcal{L}\approx\bm{z}-\bm{z}^{\prime}$.
@@
C.3.1 Supervised Training with I-MLE
The outputs of the GNN are used to steer a CO approximation algorithm. This approximation algorithm outputs a solution to the CO problem, which can be compared to the ground truth solution using a Hamming loss. During backpropagation, I-MLE (Niepert et al., 2021) is used to estimate the gradient of the loss with respect to the GNN’s output. This setting has no practical benefit over the simple supervised training using a binary cross entropy loss, but it serves to measure the effectiveness of I-MLE.
@@ C.3.2 I-MLE Target Distribution
The best value for $\theta$ is $1-y$ (we have to invert it because the input to the approximation algorithm is inverted) [^4] Using $\theta$ in the same sense as (Niepert et al., 2021). In our case, $\theta=1-s$. With this we arrive at the following target distribution:
@@
C.3.3 Self-supervised Training with I-MLE
@@
where $J_{\text{CO}}$ is the objective function of the CO problem. Note that the CO problem’s constraints don’t explicitly appear here, because the CO approximation algorithm already guarantees that the constraints are met.
As before, I-MLE is used to estimate $\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}(s-1)}$. In this setting, the general-purpose target distribution for I-MLE presented in (Niepert et al., 2021) is used.
Appendix D Graph Generation
D.1 Minimum $k$-Cut
@@
The range of possible problem instances can be improved by generating graphs of varying density. Start by assigning nodes to $k$ subgraphs of random sizes within a given range. Then, add a random number of edges that connect nodes of different subgraphs. For each subgraph, add edges between random nodes within the same subgraph until all nodes have a higher degree than the number of edges between subgraphs. As long as there are enough edges between subgraphs, the minimum $k$-cut very likely consists of the edges between subgraphs. The minimum node degree and therefore the density of the graph depends on the number of edges between subgraphs and therefore on the size of the minimum $k$-cut.
@@
The minimum $k$-cut in these graphs usually consists of the edges between subgraphs, but this is not guaranteed. The ground truth solution is therefore calculated separately to make sure that it reflects the optimal cut, as described in subsection D.3. Another benefit of calculating them separately is that we can set the number of subgraphs to a different number than $k$, generating more interesting graphs.
D.2 Travelling Salesman Problem
@@
D.3 Calculating Ground Truth Labels for Supervised Training
@@
  • Minimum k-cut: For graphs without edge weights, the graphs can be constructed with known ground truth solutions. See subsection D.1 for details. For graphs with edge weights, the Karger–Stein algorithm (Karger & Stein, 1993) is run 100 times, and the smallest cut found is treated as the ground truth minimum cut.

### v2

@@ B.2 Probability of finding optimal solutions using algorithms with relative error bounds
Appendix C Implementation details
C.1 Residual gated graph convnets
We use residual gated graph convnets (Bresson & Laurent, 2017), but adapt them to include edge features $\bm{e}_{ij}^{l}$ and a dense attention map $\bm{\eta}_{ij}^{l}$ following Joshi et al. (2019). The input node features $\bm{x}_{i}^{0}$ and edge features $\bm{e}_{ij}^{0}$ are first pre-processed using a single-layer MLP (multi-layer perceptron) for each of the two. Each further layer is computed as follows:
$\bm{x}_{i}^{l+1}$ $=\bm{x}_{i}^{l}+\mathrm{ReLU}\Bigg{(}\mathrm{BN}\bigg{(}W_{1}^{l}\bm{x}_{i}^{l}+\sum_{j\in\mathcal{N}_{i}}\bm{\eta}_{ij}^{l}\odot W_{2}^{l}\bm{x}_{j}^{l}\bigg{)}\Bigg{)}\hskip 5.69054pt\text{with }\bm{\eta}_{ij}^{l}=\frac{\sigma\big{(}\bm{e}_{ij}^{l}\big{)}}{\sum_{j^{\prime}\in\mathcal{N}_{i}}\sigma\big{(}\bm{e}_{ij^{\prime}}^{l}+\varepsilon\big{)}}\in\mathbb{R}^{d},$
The final edge-level output is calculated from the last layer’s edge features $\bm{e}_{ij}^{l}$ using another multi-layer perceptron (MLP). $f({\mathcal{G}})\in\mathbb{R}^{|E|}$ refers to applying this GNN on a graph ${\mathcal{G}}$.
C.2 Decoding using the Karger–Stein algorithm
In the case of the Karger–Stein algorithm, we noticed empirically that simply modifying the input graph can lead to degenerate behavior during testing. The Karger–Stein algorithm uses the graph’s edge weights in two places: (1) when sampling an edge for contraction and (2) when comparing the cuts that resulted from different recursion arms. We noticed that the performance of our overall method can be improved when using a model trained with the setting described in Section 5.2 by using the modified edge weights for the first case and the original edge weights for the second case. Intuitively, if the GNN makes a mistake when scaling the edge weights, using the original edge weights for comparing cuts can allow the Karger–Stein algorithm to find the optimal cut regardless.
Appendix D Details regarding baselines
D.1 Supervised training with binary cross entropy loss
The task is treated as an edge-level binary classification task. The network is trained using a BCE loss:
@@ D.1 Supervised training with binary cross entropy loss
D.2 REINFORCE
Recall that the REINFORCE algorithm (Williams, 1992), also known as the score function estimator, calculates
@@ D.2 REINFORCE
$\nabla_{\theta}J(y)\leftarrow J(y)\left(y-\displaystyle\frac{1}{N}\sum_{i=1}^{N}y_{i}\right)$
              where $\varepsilon_{i}\sim\mathrm{Gumbel}(0,1)$
                          $y_{i}\leftarrow h(\theta+\varepsilon_{i})$
@@
D.3 Implicit maximum likelihood estimator (I-MLE)
@@
In the backward pass, assume we know the gradient of the loss w.r.t. to $\bm{z}$, $\nabla_{\bm{z}}\mathcal{L}$. First, obtain a modified input $\bm{\theta}^{\prime}$ for which we can expect better outputs compared to $\bm{\theta}$. One generally applicable option suggested by Niepert et al. (2021) is $\bm{\theta}^{\prime}=\bm{\theta}-\lambda\nabla_{\bm{z}}\mathcal{L}$, where $\lambda$ is a hyperparameter. Using the same noise $\bm{\epsilon}$ as in the forward pass, perturb $\bm{\theta}^{\prime}$ and run the CO solver on $\bm{\theta}^{\prime}+\bm{\epsilon}$, obtaining $\bm{z}^{\prime}$. Finally, return the estimated gradient $\nabla_{\bm{\theta}}\mathcal{L}\approx\bm{z}-\bm{z}^{\prime}$.
@@
We used three noise samples from the Sum-of-Gamma noise distribution suggested by Niepert et al. (2021), using $\kappa=5$ and $100$ iterations.
D.3.1 Supervised training with I-MLE
The outputs of the GNN are used to guide a CO approximation algorithm. This approximation algorithm outputs a solution to the CO problem, which can be compared to the ground truth solution using a Hamming loss. During backpropagation, I-MLE (Niepert et al., 2021) is used to estimate the gradient of the loss with respect to the GNN’s output. This setting has no practical benefit over the simple supervised training using a BCE loss, but it serves to measure the effectiveness of I-MLE.
@@ D.3.2 I-MLE target distribution
The best value for $\theta$ is $1-y$ (we have to invert it because the input to the approximation algorithm is inverted). [^3] Using $\theta$ in the same sense as Niepert et al. (2021). In our case, $\theta=1-s$ With this we arrive at the following target distribution:
@@
D.3.3 Self-supervised training with I-MLE
@@
where $J_{\text{CO}}$ is the objective function of the CO problem. Note that the CO problem’s constraints do not explicitly appear here, because the CO approximation algorithm already guarantees that the constraints are met.
As before, I-MLE is used to estimate $\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}(s-1)}$. In this setting, the general-purpose target distribution for I-MLE presented in (Niepert et al., 2021) is used, setting $\lambda=20$ as suggested by that paper.
Appendix E Graph generation
E.1 Minimum $k$-cut
@@
The range of possible problem instances can be improved by generating graphs of varying density. Start by assigning nodes to $k$ subgraphs of random sizes within a given range. Then, add a random number of edges that connect nodes of different subgraphs. For each subgraph, add edges between random nodes within the same subgraph until all nodes have a higher degree than the number of edges between subgraphs. As long as there are enough edges between subgraphs, the minimum $k$-cut very likely consists of the edges between subgraphs. The minimum node degree and hence the density of the graph depends on the number of edges between subgraphs and therefore on the size of the minimum $k$-cut.
@@
The minimum $k$-cut in these graphs usually consists of the edges between subgraphs, but this is not guaranteed. The ground truth solution is therefore calculated separately to make sure that it reflects the optimal cut, as described in Section E.3. Another benefit of calculating them separately is that we can set the number of subgraphs to a different number than $k$, generating more interesting graphs.
E.2 Travelling salesman problem
@@
E.3 Calculating ground truth labels for supervised training
@@
  • Minimum k-cut: For graphs without edge weights, the graphs can be constructed with known ground truth solutions. See Section E.1 for details. For graphs with edge weights, the Karger–Stein algorithm (Karger & Stein, 1993) is run 100 times, and the smallest cut found is treated as the ground truth minimum cut.
@@
Table 3: TSP optimality gaps with mean $\pm$ standard deviation calculated over ten evaluation runs on the same model parameters. We group methods by their original paper and indicate the decoder used in parentheses. Results marked with * are values obtained from the indicated papers and therefore do not include standard deviations.
Method (decoder in parentheses) $\bm{n=20}$ $\bm{n=50}$ $\bm{n=100}$
Self-supervised
Bello et al. (2017)
   (greedy) 1.42%* 4.46%* 6.90%*
Deudon et al. (2018)
   (greedy) 0.66%* (2m) 3.98%* (5m) 8.41%* (8m)
   (greedy + 2OPT) 0.42%* (4m) 2.77%* (26m) 5.21%* (3h)
   (sampling 1280 times) 0.11%* (5m) 1.28%* (17m) 12.70%* (56m)
   (sampling 1280 times + 2OPT) 0.09%* (6m) 1.00%* (32m) 4.64%* (5h)
Khalil et al. (2017)
   (greedy) 1.42%* 5.16%* 7.03%*
Kool et al. (2019)
   (greedy) 0.34%* 1.76%* (2s) 4.53%* (6s)
   (sampling 1280 times) 0.08%* (5m) 0.52%* (24m) 2.26%* (1h)
REINFORCE (random ins., 20 runs) 7.98%$\pm$0.08 (769ms) 23.84%$\pm$0.10 (11.62s) 52.83%$\pm$0.47 (1.38m)
REINFORCE (random ins., 100 runs) 4.78%$\pm$0.03 (3.64s) 11.24%$\pm$0.08 (57.85s) 31.06%$\pm$0.25 (7.53m)
I-MLE (random ins., 20 runs) 10.54%$\pm$0.15 (761ms) 35.87%$\pm$0.34 (11.14s) 61.42%$\pm$0.23 (1.37m)
I-MLE (random ins., 100 runs) 6.55%$\pm$0.10 (3.68s) 19.40%$\pm$0.24 (56.63s) 44.42%$\pm$0.14 (7.51m)
Best-of-20 (random insertion, 20 runs) 0.40%$\pm$0.03 (787ms) 11.94%$\pm$0.11 (11.03s) 15.82%$\pm$0.09 (1.47m)
Best-of-20 (random insertion, 100 runs) 0.10%$\pm$0.01 (3.70s) 5.21%$\pm$0.15 (52.82s) 9.82%$\pm$0.09 (7.38m)
PBGE (ours) (random ins., 20 runs) 0.18%$\pm$0.01 (763ms) 2.37%$\pm$0.02 (11.11s) 5.12%$\pm$0.06 (1.43m)
PBGE (ours) (random ins., 100 runs) 0.05%$\pm$0.01 (3.73s) 1.13%$\pm$0.03 (53.98s) 3.67%$\pm$0.05 (7.44m)
Supervised, not directly comparable
Joshi et al. (2019)
   (greedy) 0.60%* (6s) 3.10%* (55s) 8.38%* (6m)
   (beam search, beam width 1280) 0.10%* (20s) 0.26%* (2m) 2.11%* (10m)
Sun & Yang (2023)
   (greedy) 0.10%* 0.24%*
   (sampling 16 times) 0.00%* 0.00%*
BCE loss (random insertion, 20 runs) 0.15%$\pm$0.01 (787ms) 0.95%$\pm$0.03 (10.92s) 2.86%$\pm$0.04 (1.47m)
BCE loss (random insertion, 100 runs) 0.04%$\pm$0.00 (3.75s) 0.59%$\pm$0.02 (54.79s) 1.75%$\pm$0.03 (7.35m)
Non-learned approximation algorithms
Random Insertion 4.46%$\pm$0.08 (41ms) 7.57%$\pm$0.08 (575ms) 9.63%$\pm$0.11 (4.34s)
LKH3 0.00%* (18s) 0.00%* (5m) 0.00%* (21m)
Exact solvers, not directly comparable
Concorde (Applegate et al., 2006) 0.00%* (1m) 0.00%* (2m) 0.00%* (3m)
Gurobi 0.00%* (7s) 0.00%* (2m) 0.00%* (17m)
Appendix F Extended comparison for TSP
Table 3 compares PBGE with more baselines, including ones that make use of supervised learning.
We also compare against a simple self-supervised baseline that runs random insertion on the input graph 20 times and treats the best solution found as ground truth for a BCE loss. We call this baseline “Best-of-20”.
The beam search decoder works similarly to the greedy decoder. It starts with an arbitrary node, then explores the $b$ edges with the highest scores. This gives us $b$ partial solutions. In each iteration, each partial solution is expanded at its last node, and out of the resulting paths, the $b$ best partial solutions are kept. Edges that would lead to invalid tours are ignored. The parameter $b$ is called the beam width, and beam search with $b=1$ corresponds to greedy search.
Appendix G Hyperparameters and other details
Table 4 and Table 5 detail the hyperparameters used for experiments on minimum $k$-cut and TSP, respectively.
Our training sets for minimum $k$-cut contain 10,000 graphs, and the validation sets contain 1,000 graphs. For TSP, our training sets consist of 1,000,000 graphs, and our validation sets contain 1,000 graphs.
Table 4: Hyperparameters used for minimum $k$-cut.
Hyperparameter Value
GNN layers 4
GNN hidden channels 32
MLP prediction head layers 2
Optimizer AdamW
Weight Decay 0.01
Learning rate scheduler ReduceLROnPlateau
Initial learning rate 0.001
Learning rate scheduler patience 4
Batch size 64 graphs
Table 5: Hyperparameters used for TSP.
Hyperparameter Value
GNN layers 12
GNN hidden channels 100
MLP prediction head layers 3
Optimizer AdamW
Weight Decay 0.01
Learning rate scheduler ReduceLROnPlateau
Initial learning rate 0.0001
Learning rate scheduler patience 4
Batch size 64 graphs
Appendix H Hardware resources used for experiments
We used an AMD EPYC 7313 16-Core processor and an NVIDIA H100 GPU in our experiments. Runtime measurements were taken on an 11th Gen Intel Core i7-11850H processor with a base speed of 2.50 GHz, using a single CPU core and no GPU, with a batch size of 1.